Which one(s) hold for any square matrix B?

You also need to explain the each correct options.Thank you2. Which one(s) hold for any square matrix B? (Herhangi bir B kare matrisi için hangileri doğrudur.) a. Determinant of B is t
Answer #1

diagonal @ (a) Let B= 1! 27 be any ZX 2 matrix Then det B= 11 2) = 1(7) - 3 (2) - 1371 = 7-6 = 1 which is not equal to producdet (8) 117-213) = 7-6 . & det B= 1 ; det 8) = det (B) (c) is alle (a) B=(8 37 [6 ] Q5783-BR, ; U = 16;] Since det B = & d